The lenght of a rectangle is one less than twice the width. If its perimeter is 34", what are the dimensions ?
Answer by htmentor(1343)

(1) 2l + 2w = 34, where l = length, w = width
(2) 2w - 1 = l, which states that the length is ONE less than TWICE the width
From (2), 2w = l + 1
Substituting into (1) gives 2l + l + 1 = 34, or 3l + 1 = 34
Solving for l gives l = 11.
Using (2), 2w - 1 = 11, or w = 12/2 = 6.
Answers: length = 11, width = 6